# Randomized Space–Time Coded Stacked Intelligent Metasurfaces for Massive Multiuser Downlink Connectivity

### v1

@@ Abstract
Stacked intelligent metasurfaces (SIMs) represent a key enabler for next-generation wireless networks, offering beamforming gains while significantly reducing radio-frequency chain requirements. In conventional space-only SIM architectures, the rate of reconfigurability of the SIM is equal to the inverse of the channel coherence time. This paper investigates a novel beamforming strategy for massive downlink connectivity using a randomized space–time (ST) coded SIM. In addition to conventional space-only metasurface layers, the proposed design integrates a ST metasurface layer at the input stage of the SIM that introduces random time variations over each channel coherence time interval. These artificial time variations enable opportunistic user scheduling and exploitation of multiuser diversity under slow channel dynamics. To mitigate the prohibitive overhead associated with full channel state information at the transmitter (CSIT), we propose a partial-CSIT-based beamforming scheme that leverages randomized steering vectors and limited user-side feedback based on signal quality measurements. Numerical results demonstrate that the proposed ST-SIM architecture achieves satisfactory sum-rate performance while significantly reducing CSIT acquisition and feedback overhead, thereby enabling scalable downlink connectivity in dense networks.
publicationid: pubid: 0000–0000/00\$00.00 © 2025 IEEE
@@ I-A Related works
The potential of SIM technology for wireless communications has been explored in several recent studies. Some works focused on point-to-point scenarios, analyzing the fundamental wave propagation mechanisms through stacked metasurfaces and the resulting beamforming capabilities. A free-space path-loss model for SIM-based transmitters was proposed in [7], while [8] introduced a multiport network model accounting for mutual coupling effects. Further developments include SIM-assisted direction-of-arrival estimation [9] and holographic multiple-input multiple-output (MIMO) systems [3], where SIMs are integrated at both the transmitter and receiver to realize parallel subchannel decomposition. In [10], the authors tackled the problem of channel estimation in SIM-assisted MIMO systems by proposing a low-overhead estimation protocol and subspace-based linear estimators that leverage the spatial correlation structure of SIM. A double-SIM-assisted massive MIMO architecture was proposed in [11], which integrates a hybrid SIM at the base station and an additional SIM in the intermediate space, jointly optimized through a projected gradient ascent method to maximize uplink spectral efficiency under imperfect channel state information (CSI).
@@
Other works addressed multiuser downlink communications with SIMs, focusing on sum-rate maximization through wave-domain beamforming. In [12], alternating optimization was used to jointly optimize transmit power and SIM transmission coefficients, while [13] employed deep reinforcement learning to address the non-convexity of the beamforming design. In [14], the authors extended the SIM-based transceiver in [12] by considering meta-atoms that can only be tuned discretely. To alleviate the high overhead associated with instantaneous CSI, [15] and [16] proposed a design based on statistical CSI, significantly simplifying system operation in slowly varying channels. In [17], the authors proposed a hybrid transceiver architecture for near-field wideband systems assisted by SIM, introducing a layer-by-layer holographic beamforming algorithm combined with minimum-mean-square-error digital precoding to maximize spectral efficiency under realistic phase tuning errors. All these studies considered phase-only SIM, which are nearly passive and easier to implement but suffer from uncontrollable propagation losses across multiple layers. A more recent contribution has investigated the inclusion of active amplitude-controlling layers to enhance wave manipulation capabilities and mitigate such internal losses [18].
@@
Recently, the integration of SIM into cell-free massive MIMO architectures has been studied to address scalability and fronthaul bottlenecks. A digital-wave beamforming framework was introduced in [19], where SIM-equipped access points enable high beamforming gains with fewer RF chains. Alternating optimization algorithms were developed in [20] for the joint design of digital and wave-domain beamforming, as well as fronthaul compression, showing that sufficiently deep SIMs can approach the performance of fully digital schemes. Other works [21, 22, 23] demonstrated SIMs’ ability to improve both uplink and downlink sum-rate performance while substantially reducing hardware and fronthaul costs.
@@ I-B Contributions
    We introduce a novel randomized space–time (ST) coded SIM architecture that incorporates a rapidly time-varying (TV) dimensional adaptation layer, which is randomly reconfigured at a rate greater than $1/T$, and multiple space-only metasurface layers, whose characteristic parameters vary slowly at a rate of $1/T$. This design enables joint spatial-temporal wavefront control, introducing artificial time variations over each channel coherence time interval that enhance multiuser diversity even under slowly varying propagation conditions.
@@ I-C Paper organization
The remainder of this paper is organized as follows. Section II introduces the system model and problem formulation for the proposed ST-SIM architecture. Section III outlines the mathematical models of both transmitted and received signals. Section IV presents the partial-CSI scheduling strategy. In Section V, the optimization framework for designing the SIM transmission coefficients is developed relying on the gradient descent algorithm. Section VI provides illustrates results and performance comparisons with conventional beamforming schemes. Finally, Section VII provides some concluding remarks.
@@
Figure 1: ST coded SIM-aided multiuser downlink system serving NN out of $U$ users. The ST-SIM consists of $L$ metasurface layers. The first layer acts as a ST-coded DAL, comprising both absorbing (in red color) and transmitting (in green color) meta-atoms. The last L1L-1 layers are S-coded ones and consist only of transmitting meta-atoms. The S-coded layers do not vary over each channel coherence interval of duration $T$, while the ST-coded initial layer is reconfigured at a rate $M$ times larger than $1/T$.
II Space-time coded SIM architecture
@@
We now introduce the system model underpinning the proposed ST-SIM architecture. The baseband modulation spectrum is confined to a frequency interval $\mathcal{W}_{\mathrm{m}}$ of width $B_{\mathrm{m}}$ (bandwidth) centered at $f=0$, with $B_{\mathrm{m}}\ll f_{0}$ (narrowband assumption), where $f_{0}$ denotes the carrier frequency. Let $T$ denote the channel coherence time, which is inversely related to the bandwidth of the Doppler spectrum of the underlying physical channel. We consider a transmission interval $[0,T)$, which is partitioned into $M$ time slots of duration $T_{\text{s}}$, i.e., $T=M\,T_{\text{s}}$. We utilize a uniform planar array (UPA) to radiate the information streams, which consists of $N\triangleq N_{x}\times N_{y}$ transmit antennas arranged in a rectangular grid with $N_{x}$ and $N_{y}$ elements along the $x$ and $y$ axes, respectively, and inter-element spacing $d_{\text{upa}}$. As depicted in Fig. 1, the base station (BS) is composed of the UPA and the ST-SIM. [^1] In Fig. 1, the UPA is depicted as a linear array of $N$ antennas with inter-element spacing $d_{\text{upa}}$. Such a ST-SIM-based transmitter operates on a slot-by-slot basis.
@@
The ST-coded SIM comprises $L$ planar layers, uniformly spaced by a distance $s_{\text{lay}}$ (see Fig. 1). Specifically, it consists of two functional blocks with different rate of reconfigurability: (i) a ST-coded block (in light-purple color in Fig. 1), implemented by the initial metasurface layer with transmission coefficients updated at every time slot, i.e., the transmission properties of its meta-atoms are reconfigured at rate $f_{\text{s}}\triangleq 1/T_{\text{s}}$ (rapidly TV); (ii) a space-only-coded (S-coded) block (in light-blue color in Fig. 1), comprising the subsequent L1L-1 metasurface layers with transmission coefficients fixed over each transmission interval of duration $T$, i.e., the response of such layers is reconfigured at rate $1/T$ (slowly TV). It is noteworthy that the rate of reconfigurability of the first ST-coded layer of the SIM is $M$ times larger than that of the other S-coded layers. This two-timescale layered architecture enables joint spatial and temporal wavefront processing across the metasurface stack. The transmitting meta-atoms of the TV layer are time-modulated under the adiabatic condition $f_{\text{s}}\ll f_{0}$ [24].
@@
Regarding the meta-atom type, the $L$ metasurface layers of the SIM can be further partitioned into two groups: the boundary layers ($1$ and $L$) and the L2L-2 intermediate layers ($2,\ldots,L-1$). Each of the intermediate L2L-2 layers consists of $Q\triangleq Q_{x}\times Q_{y}$ transmitting meta-atoms (in green color in Fig. 1) arranged in a rectangular grid, with $Q_{x}$ and $Q_{y}$ elements along the $x$- and $y$-axes, respectively, and inter-element spacing $d_{\text{meta}}$. Hereinafter, the inter-element spacing of each metasurface layer $d_{\text{meta}}$ is assumed, for simplicity, to be equal to the inter-antenna spacing of the UPA $d_{\text{upa}}$, i.e., $d_{\text{upa}}=d_{\text{meta}}$. By contrast, the boundary layers contain fewer than $Q$ transmitting meta-atoms. The first ST-coded layer comprises $Z\triangleq Z_{x}\times Z_{y}\leq Q$ transmitting meta-atoms spatially aligned with the UPA grid ($Z\geq N$), with inter-element spacing $d_{\text{meta}}$; the remaining QZQ-Z meta-atoms are perfectly absorbing and surround the transmitting region (in red color in Fig. 1). Likewise, the terminal layer $L$ comprises $V\triangleq V_{x}\times V_{y}\leq Q$ transmitting meta-atoms aligned with the same grid ($V\geq N$), with spacing $d_{\text{meta}}$, while the remaining QVQ-V elements are perfectly absorbing. These layers are designed to decouple the dimensionality of the ST-SIM from the number of transmitting meta-atoms $Q$. To highlight this property, we refer to the first and last metasurfaces as dimensional adaptation layers (DALs).
@@
In a conventional SIM architecture with dimensions $Q\times N$, the number of variables to be optimized is $QN$, while the degrees of freedom at the designer’s disposal are limited to the number of metasurface layers $L$. Consequently, the designer is constrained to increasing $L$ to enhance performance, which may adversely affect the convergence rate of the iterative algorithms employed to synthesize the transmission SIM coefficients, due to error propagation across layers. In contrast, the proposed SIM architecture decouples the number of transmitting antennas $N$ and the number of meta-atoms $Q$ of the intermediate L2L-2 layers from the number of elements of the overall SIM response matrix, having dimensions $V\times Z$, thereby introducing $Q$ as an independent design variable that can be optimized as the number of layers $L$. It is worth noting that the inclusion of the DAL is applicable to any SIM design, which is not limited to ST-coded implementations only.
@@
When an incident EM wave impinges on a meta-atom, the transmitted wave’s amplitude and phase are determined by the product of the incident field and the meta-atom’s complex-valued transmission coefficient. The re-radiated wave then serves as a secondary source illuminating the subsequent layer, following the Huygens–Fresnel principle [25]. On the contrary, each wave striking one of the absorbing meta-atoms in the DALs is completely dissipated by integrated circuitry, preventing further propagation. From a hardware perspective, a strongly attenuating (almost absorbing) response can be realized using purely passive lossy meta-atoms. In our architecture, however, some layers already employ active meta-atoms with integrated amplifier chips to enable programmable amplitude control. Under field programmable gate array (FPGA) control, these meta-atoms operate active artificial neurons, enabling a large dynamic modulation range (e.g., from 22-22 to $13$ dB in [5]). Therefore, these same active elements can also be driven into a highly lossy operating point, where they effectively behave as near-perfect absorbers, strongly suppressing the propagation of the incident wave to subsequent layers. Hereinafter, for modeling convenience, the meta-atoms in the DALs will be treated as perfect absorbers, which fully dissipate the incident energy.

### v2

# Randomized Space-Time Stacked Intelligent Metasurfaces for Massive Multiuser Downlink Connectivity

@@ Abstract
Stacked intelligent metasurfaces (SIMs) represent a key enabler for next-generation wireless networks, offering beamforming gains while significantly reducing radio-frequency chain requirements. In conventional space-only SIM architectures, the rate of reconfigurability of the SIM is equal to the inverse of the channel coherence time. This paper investigates a novel beamforming strategy for massive downlink connectivity using a randomized space–time (ST) SIM. In addition to conventional space-only metasurface layers, the proposed design integrates a ST metasurface layer at the input stage of the SIM that introduces random time variations over each channel coherence time interval. These artificial time variations enable opportunistic user scheduling and exploitation of multiuser diversity under slow channel dynamics. To mitigate the prohibitive overhead associated with full channel state information at the transmitter (CSIT), we propose a partial-CSIT-based beamforming scheme that leverages randomized steering vectors and limited user-side feedback based on signal quality measurements. Numerical results demonstrate that the proposed ST-SIM architecture achieves satisfactory sum-rate performance while significantly reducing CSIT acquisition and feedback overhead, thereby enabling scalable downlink connectivity in dense networks.
publicationid: pubid: 0000–0000/00\$00.00 © 2026 IEEE
@@ I-A Related works
The potential of SIM technology for wireless communications has been explored in several recent studies. Some works focused on point-to-point scenarios, analyzing the fundamental wave propagation mechanisms through stacked metasurfaces and the resulting beamforming capabilities. A free-space path-loss model for SIM-based transmitters was proposed in [7], while [8] introduced a multiport network model accounting for mutual coupling effects. Further developments include SIM-assisted direction-of-arrival estimation [9] and holographic multiple-input multiple-output (MIMO) systems [3], where SIMs are integrated at both the transmitter and receiver to realize parallel subchannel decomposition. In [10], the authors tackled the problem of channel estimation in SIM-assisted MIMO systems by proposing a low-overhead estimation protocol and subspace-based linear estimators that leverage the SIM spatial correlation structure. A double-SIM-assisted massive MIMO architecture was proposed in [11], which integrates a hybrid SIM at the base station and an additional SIM in the intermediate space, jointly optimized through a projected gradient ascent method to maximize uplink spectral efficiency under imperfect channel state information (CSI).
@@
Other works addressed multiuser downlink communications with SIMs, focusing on sum-rate maximization through wave-domain beamforming. In [12], alternating optimization was used to jointly optimize transmit power and SIM transmission coefficients, while [13] employed deep reinforcement learning to address the non-convexity of the beamforming design. In [14], the authors extended the SIM-based transceiver in [12] by considering meta-atoms that can only be tuned discretely. To alleviate the high overhead associated with instantaneous CSI, [15] and [16] proposed a design based on statistical CSI, significantly simplifying system operation in slowly varying channels. In [17], the authors proposed a hybrid transceiver architecture for near-field wideband systems assisted by SIM, introducing a layer-by-layer holographic beamforming algorithm combined with minimum-mean-square-error digital precoding to maximize spectral efficiency under realistic phase tuning errors. All these studies considered phase-only SIMs, which are nearly passive and easier to implement but suffer from uncontrollable propagation losses across multiple layers. A more recent contribution has investigated the inclusion of active amplitude-controlling layers to enhance wave manipulation capabilities and mitigate such internal losses [18].
@@
Recently, the integration of SIMs into cell-free massive MIMO architectures has been studied to address scalability and fronthaul bottlenecks. A digital-wave beamforming framework was introduced in [19], where SIM-equipped access points enable high beamforming gains with fewer RF chains. Alternating optimization algorithms were developed in [20] for the joint design of digital and wave-domain beamforming, as well as fronthaul compression, showing that sufficiently deep SIMs can approach the performance of fully digital schemes. Other works [21, 22, 23] demonstrated SIMs’ ability to improve both uplink and downlink sum-rate performance while substantially reducing hardware and fronthaul costs.
Recent studies have also investigated SIMs for enhancing the spectral and energy efficiency of multiuser systems. In particular, the authors in [24] analyze the energy efficiency of SIM-assisted communications by jointly optimizing digital precoding and wave-based beamforming, with special emphasis on the impact of the number of metasurface layers under realistic power consumption models.
Hardware nonidealities constitute an important practical aspect in holographic and SIM-based wireless architectures, including phase and amplitude imperfections at metasurface elements as well as impairments in the associated RF chains. The impact of such nonidealities has been investigated for SIM-assisted cell-free networks in [25], while impairment-aware holographic metasurface beamforming has been analyzed for multi-altitude low earth orbit (LEO) satellite systems in [26]. Moreover, the performance of reconfigurable holographic surfaces in near-field cell-free scenarios under phase-shift errors and RF impairments has been theoretically characterized in [27].
@@ I-B Contributions
    We introduce a novel randomized space–time (ST) SIM architecture that incorporates a rapidly time-varying (TV) dimensional adaptation layer, which is randomly reconfigured at a rate greater than $1/T$, and multiple space-only metasurface layers, whose characteristic parameters vary slowly at a rate of $1/T$. This design enables joint spatial-temporal wavefront control, introducing artificial time variations over each channel coherence time interval that enhance multiuser diversity even under slowly varying propagation conditions.
@@
It is worth noting that the proposed ST-SIM should not be confused with the space–time coding metasurfaces introduced in [28], which exploit periodic, phase-quantized time modulation to engineer harmonic spectra and control frequency conversion. In contrast, the ST layer considered in this work introduces randomized phase variations across time slots in order to induce artificial channel fluctuations and enable opportunistic multiuser scheduling.
@@ I-C Paper organization
The remainder of this paper is organized as follows. Section II introduces the system model and problem formulation for the proposed ST-SIM architecture. Section III outlines the mathematical models of both transmitted and received signals. Section IV presents the partial-CSI scheduling strategy. In Section V, the optimization framework for designing the SIM transmission coefficients is developed relying on the gradient descent algorithm. Section VI provides illustrative numerical results and performance comparisons with conventional beamforming schemes. Finally, Section VII provides some concluding remarks.
@@
Figure 1: ST-SIM-aided multiuser downlink system serving NN out of $U$ users. The ST-SIM consists of $L$ metasurface layers. The first layer acts as a ST-DAL, comprising both absorbing (in red color) and transmitting (in green color) meta-atoms. The last L1L-1 layers are S-only ones and consist of transmitting meta-atoms. S-only layers do not vary over each channel coherence interval of duration $T$, while the ST initial layer is reconfigured at a rate $M/T$.
II Space-time SIM architecture
@@
Herein, we introduce the system model underpinning the proposed ST-SIM architecture. The baseband modulation spectrum is confined to a frequency interval $\mathcal{W}_{\mathrm{m}}$ of width $B_{\mathrm{m}}$ (bandwidth) centered at $f=0$, with $B_{\mathrm{m}}\ll f_{0}$ (narrowband assumption), where $f_{0}$ denotes the carrier frequency. Let $T$ denote the channel coherence time, which is inversely related to the bandwidth of the Doppler spectrum of the underlying physical channel. We consider a transmission interval $[0,T)$, which is partitioned into $M$ time slots of duration $T_{\text{s}}$, i.e., $T=M\,T_{\text{s}}$. We utilize a uniform planar array (UPA) to radiate the information streams, which consists of $N\triangleq N_{x}\times N_{y}$ transmit antennas arranged in a rectangular grid with $N_{x}$ and $N_{y}$ elements along the $x$ and $y$ axes, respectively, and inter-element spacing $d_{\text{upa}}$. As depicted in Fig. 1, the base station (BS) is composed of the UPA and the ST-SIM. [^1] In Fig. 1, the UPA is depicted as a linear array of $N$ antennas with inter-element spacing $d_{\text{upa}}$. Such a ST-SIM-based transmitter operates on a slot-by-slot basis.
@@
The ST-SIM comprises $L$ planar layers, uniformly spaced by a distance $s_{\text{lay}}$ (see Fig. 1). Specifically, it consists of two functional blocks with different rate of reconfigurability: (i) a ST block (in light-purple color in Fig. 1), implemented by the initial metasurface layer with transmission coefficients updated at every time slot, i.e., the transmission properties of its meta-atoms are reconfigured at rate $f_{\text{s}}\triangleq 1/T_{\text{s}}$ (rapidly TV); (ii) a space-only (S-only) block (in light-blue color in Fig. 1), comprising the subsequent L1L-1 metasurface layers with transmission coefficients fixed over each transmission interval of duration $T$, i.e., the response of such layers is reconfigured at rate $1/T$ (slowly TV). It is noteworthy that the rate of reconfigurability of the first ST layer of the SIM is $M$ times larger than that of the other S-only layers. This two-timescale layered architecture enables joint spatial and temporal wavefront processing across the metasurface stack. The transmitting meta-atoms of the TV layer are time-modulated under the adiabatic condition $f_{\text{s}}\ll f_{0}$ [29]. In contrast to [25, 26, 27], the present work focuses on ST-SIM architectures for inducing artificial channel fluctuations and enabling opportunistic multiuser scheduling under partial CSIT, while assuming ideal hardware in order to isolate the fundamental wave-domain effects.
@@
Regarding the meta-atom type, the $L$ metasurface layers of the SIM can be further partitioned into two groups: the boundary layers ($1$ and $L$) and the L2L-2 intermediate layers ($2,\ldots,L-1$). Each of the intermediate L2L-2 layers consists of $Q\triangleq Q_{x}\times Q_{y}$ transmitting meta-atoms (in green color in Fig. 1) arranged in a rectangular grid, with $Q_{x}$ and $Q_{y}$ elements along the $x$- and $y$-axes, respectively, and inter-element spacing $d_{\text{meta}}$. Hereinafter, the inter-element spacing of each metasurface layer $d_{\text{meta}}$ is assumed, for simplicity, to be equal to the inter-antenna spacing of the UPA $d_{\text{upa}}$, i.e., $d_{\text{upa}}=d_{\text{meta}}$. By contrast, the boundary layers contain fewer than $Q$ transmitting meta-atoms. The first ST layer comprises $Z\triangleq Z_{x}\times Z_{y}\leq Q$ transmitting meta-atoms spatially aligned with the UPA grid ($Z\geq N$), with inter-element spacing $d_{\text{meta}}$; the remaining QZQ-Z meta-atoms are perfectly absorbing and surround the transmitting region (in red color in Fig. 1). Likewise, the terminal layer $L$ comprises $V\triangleq V_{x}\times V_{y}\leq Q$ transmitting meta-atoms aligned with the same grid ($V\geq N$), with spacing $d_{\text{meta}}$, while the remaining QVQ-V elements are perfectly absorbing. These layers are designed to decouple the dimensionality of the ST-SIM from the number $Q$ of transmitting meta-atoms. In order to highlight this property, we refer to the first and last metasurfaces as dimensional adaptation layers (DALs).
TABLE I: Comparison between a conventional SIM and the proposed DAL-aided ST-SIM.
Characteristic Conventional SIM Proposed DAL-aided ST-SIM
SIM modulation Space-only (fixed over the coherence interval $T$) Space–time (ST layer varies across time slots of duration $T_{s}$, with $T=MT_{s}$)
Boundary layers All transmitting meta-atoms DALs include transmitting and absorbing meta-atoms: the first ST-DAL has $Z$ transmitting and QZQ-Z absorbing meta-atoms, while the last S-only DAL has $V$ transmitting and QVQ-V absorbing meta-atoms
Intermediate layers $Q$ transmitting meta-atoms Same: $Q$ transmitting meta-atoms
Response dimension Tied to $Q$: ${\mathbf{G}}\in\mathbb{C}^{Q\times N}$ Decoupled from $Q$: ${\mathbf{G}}(t)\in\mathbb{C}^{V\times N}$
Degrees of control Mainly via $L$ Via both $L$ and $Q$
@@
In a conventional SIM architecture with dimensions $Q\times N$, the number of variables to be optimized is $QN$, while the degrees of freedom at the designer’s disposal are limited to the number of metasurface layers $L$. Consequently, the designer is constrained to increasing $L$ to enhance performance, which may adversely affect the convergence rate of the iterative algorithms employed to synthesize the transmission SIM coefficients, due to error propagation across layers. In contrast, the proposed SIM architecture decouples the number of transmitting antennas $N$ and the number of meta-atoms $Q$ of the intermediate L2L-2 layers from the number of elements of the overall SIM response matrix, having dimensions $V\times Z$, thereby introducing $Q$ as an independent design variable that can be optimized as the number of layers $L$. It is worth noting that the inclusion of the DAL is applicable to any SIM design, which is not limited to ST implementations only.
@@
When an incident EM wave impinges on a meta-atom, the transmitted wave’s amplitude and phase are determined by the product of the incident field and the meta-atom’s complex-valued transmission coefficient. The re-radiated wave then serves as a secondary source illuminating the subsequent layer, following the Huygens–Fresnel principle [30]. On the contrary, each wave striking one of the absorbing meta-atoms in the DALs is completely dissipated by integrated circuitry, preventing further propagation. From a hardware perspective, a strongly attenuating (almost absorbing) response can be realized using purely passive lossy meta-atoms. In our architecture, however, some layers already employ active meta-atoms with integrated amplifier chips to enable programmable amplitude control. Under field programmable gate array (FPGA) control, these meta-atoms operate active artificial neurons, enabling a large dynamic modulation range (e.g., from 22-22 to $13$ dB in [5]). Therefore, these same active elements can also be driven into a highly lossy operating point, where they effectively behave as near-perfect absorbers, strongly suppressing the propagation of the incident wave to subsequent layers. While perfect absorption over a wide bandwidth is difficult to achieve [31, 32], moderate deviations from ideal absorption primarily result in small reflections that can be interpreted as perturbations of the effective transmission coefficients of the metasurface layers. Such effects do not alter the operation of the proposed synthesis algorithm, which optimizes the metasurface transmission coefficients to approximate the desired transfer matrix. Hereinafter, for modeling convenience, the meta-atoms in the DALs will be treated as perfect absorbers, which fully dissipate the incident energy.
Table I summarizes the key structural differences with respect to conventional SIM architectures.